# MeAJOR Corpus: A Multi-Source Dataset for Phishing Email Detection

### v1

@@ IV Baseline Experiments
Analysing the results, it can be noticed that the RF model showed significant improvement when URL features were included alongside textual data, achieving an F1 of 96.63%. This corroborates findings by Naswir et al. [49] regarding the value of URL features, though our implementation uses a more limited feature set than their hybrid approaches. In contrast, attachment features showed a neutral or slightly negative impact on the overall performance, consistent with Zhang et al. [56] who observed minimal contribution from attachments in their large-scale study. Comparing the Text + Attachment and Text + URL + Attachment feature sets with their counterparts without the attachment features, confirms that URL information is particularly valuable for phishing detection when using RF, while attachment information has little to no impact.
XGB consistently delivered strong results across all feature combinations, achieving 98.34% F1 with the Text + URL Set. This exceeds Omotehinwa et al. [48] (98.16% F1) and approaches ensemble methods like Chanis and Arampatzis [50] (98.43% F1), despite our simpler feature engineering. The robustness of the Text Only Set (97.89% F1) supports Khalid et al. [11] on text-centric efficacy, though our URL augmentation shows greater gains than their stylometric features. Attachment features again demonstrated neutral-to-negative impacts.
The MLP also exhibited strong results across all feature sets, reaching its best performance at 97.69% F1 with the Text + Attachment Set, validating Altwaijry et al. [12] on neural networks’ text proficiency. Notably, URLs caused a 1.1% F1 drop, contrasting sharply with ensemble models and highlighting MLP’s ineffectiveness with structured features (a limitation not observed in hybrid architectures like the one from Atawneh and Aljehani [51]).
Finally, the CNN model, evaluated on the same feature sets, achieved its best result when combining textual and URL data, reaching 97.66% F1 with the Text + URL Set. This demonstrated superior structural feature handling versus MLP, though falling short of Kaushik and Rathore’s work [52] (99% F1), where a specialised URL-to-image transformation is used to capture intricate visual patterns within URLs. The CNN model also maintained high precision and recall across configurations, showing its ability to generalise well even on diverse feature inputs. Notably, adding attachment features showed slightly more positive impacts when compared to the ensemble models, suggesting CNN’s adaptability to heterogeneous inputs aligns with Zhang et al. [55] on multi-modal processing.
@@ V Discussion
The results obtained from the baseline experiments are crucial for validating the effectiveness, representativeness, and utility of the newly generated dataset presented in this study. Our results robustly confirm the established trend, highlighted by Alhuzali et al. [7] and Chanis et al. [50], that ensemble models such as XGB and RF consistently excel with structured email features. XGB emerged as the top performer in our evaluation, achieving a peak F1 of 98.34% using the combined Text + URL Set. This performance is highly competitive with recent state-of-the-art results, as it surpasses the 98.16% F1 reported by Omotehinwa et al. [48] using XGB on the Enron-Spam corpus and approaches the 98.43% F1 achieved by Chanis et al. [50] using a sophisticated stacked ensemble incorporating stylometric features. Importantly, our result was obtained with comparatively simpler feature engineering, underscoring the quality and richness of the underlying dataset. The strong performance of the Text Only Set (97.89% F1) aligns with findings by Khalid et al. [11] on the efficacy of text-centric approaches. However, our integration of URL features presented more substantial gains than the stylometric augmentations reported in their LogiTriBlend model.
@@
The DL models, CNN and MLP, also demonstrated robust performance, particularly with enriched feature sets, supporting findings by Atawneh et al. [51] and Kaushik et al. [52] on the applicability of DL to textual features. However, our results reveal the interesting nuance that classical MLP architectures matched or even slightly surpassed CNN performance in some configurations, especially when incorporating structured features like attachments. This contrasts with some state-of-the-art hybrid DL approaches [51, 52], but highlights a potential advantage of simpler neural architectures when robust feature engineering provides discriminative inputs rather than relying solely on end-to-end representation learning. Our CNN model’s peak F1 of 97.66% with the Text + URL Set falls short of Kaushik et al. 99% [52], which leveraged specialised URL-to-image transformations. This gap suggests potential for future work incorporating similar advanced feature representations using our dataset.
@@
The significant performance boost observed when adding URL features aligns strongly with the current state of the art, reinforcing findings by Khalid et al. [11], Naswir et al. [49], and Zhang et al. [55] on the critical importance of lexical and structural URL characteristics for identifying phishing attempts. While attachment features showed a more neutral or context-dependent impact in our experiments (consistent with the minimal contribution noted by Zhang et al. [56]), their inclusion reflects the multi-modal learning trend emphasised in cutting-edge research [56, 41]. Our dataset’s design inherently supports this multi-modal approach by providing diverse feature categories.
@@ Acknowledgment
This work has been supported by the PC2phish project, which has received funding from FCT with Refª: 2024.07648.IACDC. Furthermore, this work also received funding from the project UIDB/00760/2020.

### v2

@@ IV Baseline Experiments
Analysing the results, it can be noticed that the RF model showed significant improvement when URL features were included alongside textual data, achieving an F1 of 96.63%. This corroborates findings by Naswir et al. [Naswir2022] regarding the value of URL features, though our implementation uses a more limited feature set than their hybrid approaches. In contrast, attachment features showed a neutral or slightly negative impact on the overall performance, consistent with Zhang et al. [Zhang2025] who observed minimal contribution from attachments in their large-scale study. Comparing the Text + Attachment and Text + URL + Attachment feature sets with their counterparts without the attachment features, confirms that URL information is particularly valuable for phishing detection when using RF, while attachment information has little to no impact.
XGB consistently delivered strong results across all feature combinations, achieving 98.34% F1 with the Text + URL Set. This exceeds Omotehinwa et al. [Omotehinwa2023] (98.16% F1) and approaches ensemble methods like Chanis and Arampatzis [Chanis2024] (98.43% F1), despite our simpler feature engineering. The robustness of the Text Only Set (97.89% F1) supports Khalid et al. [Khalid2024] on text-centric efficacy, though our URL augmentation shows greater gains than their stylometric features. Attachment features again demonstrated neutral-to-negative impacts.
The MLP also exhibited strong results across all feature sets, reaching its best performance at 97.69% F1 with the Text + Attachment Set, validating Altwaijry et al. [Altwaijry2024] on neural networks’ text proficiency. Notably, URLs caused a 1.1% F1 drop, contrasting sharply with ensemble models and highlighting MLP’s ineffectiveness with structured features (a limitation not observed in hybrid architectures like the one from Atawneh and Aljehani [Atawneh2023]).
Finally, the CNN model, evaluated on the same feature sets, achieved its best result when combining textual and URL data, reaching 97.66% F1 with the Text + URL Set. This demonstrated superior structural feature handling versus MLP, though falling short of Kaushik and Rathore’s work [Kaushik2023] (99% F1), where a specialised URL-to-image transformation is used to capture intricate visual patterns within URLs. The CNN model also maintained high precision and recall across configurations, showing its ability to generalise well even on diverse feature inputs. Notably, adding attachment features showed slightly more positive impacts when compared to the ensemble models, suggesting CNN’s adaptability to heterogeneous inputs aligns with Zhang et al. [Zhang2016] on multi-modal processing.
@@ V Discussion
The results obtained from the baseline experiments are crucial for validating the effectiveness, representativeness, and utility of the newly generated dataset presented in this study. Our results robustly confirm the established trend, highlighted by Alhuzali et al. [Alhuzali2025] and Chanis et al. [Chanis2024], that ensemble models such as XGB and RF consistently excel with structured email features. XGB emerged as the top performer in our evaluation, achieving a peak F1 of 98.34% using the combined Text + URL Set. This performance is highly competitive with recent state-of-the-art results, as it surpasses the 98.16% F1 reported by Omotehinwa et al. [Omotehinwa2023] using XGB on the Enron-Spam corpus and approaches the 98.43% F1 achieved by Chanis et al. [Chanis2024] using a sophisticated stacked ensemble incorporating stylometric features. Importantly, our result was obtained with comparatively simpler feature engineering, underscoring the quality and richness of the underlying dataset. The strong performance of the Text Only Set (97.89% F1) aligns with findings by Khalid et al. [Khalid2024] on the efficacy of text-centric approaches. However, our integration of URL features presented more substantial gains than the stylometric augmentations reported in their LogiTriBlend model.
@@
The DL models, CNN and MLP, also demonstrated robust performance, particularly with enriched feature sets, supporting findings by Atawneh et al. [Atawneh2023] and Kaushik et al. [Kaushik2023] on the applicability of DL to textual features. However, our results reveal the interesting nuance that classical MLP architectures matched or even slightly surpassed CNN performance in some configurations, especially when incorporating structured features like attachments. This contrasts with some state-of-the-art hybrid DL approaches [Atawneh2023, Kaushik2023], but highlights a potential advantage of simpler neural architectures when robust feature engineering provides discriminative inputs rather than relying solely on end-to-end representation learning. Our CNN model’s peak F1 of 97.66% with the Text + URL Set falls short of Kaushik et al. 99% [Kaushik2023], which leveraged specialised URL-to-image transformations. This gap suggests potential for future work incorporating similar advanced feature representations using our dataset.
@@
The significant performance boost observed when adding URL features aligns strongly with the current state of the art, reinforcing findings by Khalid et al. [Khalid2024], Naswir et al. [Naswir2022], and Zhang et al. [Zhang2016] on the critical importance of lexical and structural URL characteristics for identifying phishing attempts. While attachment features showed a more neutral or context-dependent impact in our experiments (consistent with the minimal contribution noted by Zhang et al. [Zhang2025]), their inclusion reflects the multi-modal learning trend emphasised in cutting-edge research [Zhang2025, Yoon2024]. Our dataset’s design inherently supports this multi-modal approach by providing diverse feature categories.
@@ Acknowledgment
This research was funded by the Itea Project VESTA (#21011) and Project VESTA (NORTE2030-FEDER-00382400) co-funded by Portugal 2030. Furthermore, this work also received funding from the project UIDB/00760/2020.